# AttackSeqBench: Benchmarking Large Language Models in Analyzing Attack Sequences within Cyber Threat Intelligence

### v2

@@ 3.1 Benchmark Settings
Zero-shot Setting. Motivated by the significant zero-shot reasoning abilities of LLMs across several downstream tasks (Hou et al., 2024; Kojima et al., 2022), we directly utilize the system prompt and the Q&A pairs to evaluate the LLMs’ performance on three tasks based on their inherent knowledge.
Context Setting. Considering the existing context-aware work (Ma et al., 2023; Jin et al., 2024), we also organize this context setting to evaluate LLMs’ Domain-Specific Epistemic Scalability at the semantic level. Here, we remove the corresponding summary of the ground truth tactic $t_{k}$ from the CTI outline $O$ to construct the masked CTI outline $O_{m}$, where $O_{m}\!=\!O\setminus o_{k}$. Afterwards, the LLM will be instructed to use the corresponding $O_{m}$ to answer the question, highlighting its potential to perform abductive reasoning to determine the most plausible TTP in attack sequence.
RAG-empowered Setting. Previous studies have demonstrated the Retrieval Augmented Generation (RAG) can significantly enhance the reliability pf LLMs and mitigate hallucinations (Zhang et al., 2025a). Here, we also design the RAG-empowered setting to evaluate the Domain-Specific Epistemic Scalability of LLMs at the representation level. This leverages the LLMs’ in-context learning ability to learn the associations between the entity in the question’s body and the relevant TTPs, thereby decomposing the problem and eliciting its stronger reasoning ability (Wu et al., 2022).
3.2 Implementation Details
To investigate the CRU capability of existing models, we evaluate seven LLMs (i.e., LLaMa-3.1-8B (Grattafiori et al., 2024), ChatGLM-4-9B (GLM et al., 2024), Qwen-2.5-3B, Qwen-2.5-14B, Qwen-2.5-32B (QwenTeam, 2024), Llama-3.3-70B (Grattafiori et al., 2024), and GPT-4o (OpenAI, 2024a) ) and five LRMs (i.e., R1 (Llama-3.1-8B), R1(Qwen-2.5-14B), R1(Qwen-2.5-32B) (DeepSeek-AI, 2025), QWQ-32B (Team, 2024) and GPT-o3-mini (OpenAI, 2025)) on AttackSeqBench. We also utilize four post-training strategies (i.e., SFT (Zhang et al., 2023), RD (Huang et al., 2024), RLIF (Zhao et al., 2025) and RLVR (DeepSeek-AI, 2025)) to embed security knowledge into LLMs to evaluate the Domain-Specific Epistemic Scalability of our AttackSeqBench (cf. Appendix A.4). Here, we measure the performance with accuracy Acc=n/MAcc\!=\!n/M, where $n$ is the correctly-answered number of questions and $M$ is the total number.  [^1] Due to the page limitation, we introduce the complete implementation details in Appendix A.5 and A.6.
Table 2: Performance comparison of various LLMs, LRMs and post-training strategies across three benchmark tasks and settings. Bold and underlined denote best and second-best in each column.
LLMs AttackSeq-Tactic AttackSeq-Technique AttackSeq-Procedure
@@ 3.2 Implementation Details
3.3 Performance Comparison
Comparison between diverse groups of LLMs. As shown in Table 2, we notice that: Although LLMs generally follow the scaling laws in our AttackSeqBench, none of the LLMs consistently outperforms the others, and the optimal LLM varies diverse tasks. For instance, the best-performing models under the zero-shot setting across the three benchmark tasks are Qwen-2.5-32B, Qwen-2.5-32B, and GPT-o3-mini, respectively. This suggests that current models may not possess explicit security-specific knowledge, as relevant information in pretraining corpus is likely overshadowed by general-domain content. Moreover, most models consistently perform worst in AttackSeq-Tactic compared to the other two tasks, mirroring the human evaluation results in Section 2.3 and underscoring the common challenge faced by both human experts and LLMs in tactical inference.
Furthermore, from Table 3, we can observe that compared to the zero-shot setting, all models exhibit substantial performance gains on AttackSeq-Procedure-No under the context setting, indicating the importance of contextual information in identifying highly implausible actions within attack sequences. As defined in Appendix A.3, AttackSeq-Procedure-No is inherently more complex and reasoning-demanding than AttackSeq-Procedure-Yes, as it requires models to overcome the helpful-only bias and explicitly answer ‘No’ to disprove the plausibility of procedures occurring within the attack sequence. This explains why LRMs with stronger reasoning ability outperform in AttackSeq-Procedure-No compared to other tasks, underscoring the benchmark’s emphasis on Reasoning Scalability. Finally, most post-training strategies substantially improve the performance of its base LLM, particularly in zero-shot settings that rely solely on internal knowledge. However, their performance still lags behind instructive LLMs equipped with task-adapted prompts. This highlights a promising direction: Designing specialized post-training strategy to embed security-related knowledge into existing LLMs, thereby advancing the development of domain-specific models for cybersecurity.
Table 3: Performance comparison of two LLMs and two LRMs on AttackSeq-Procedure-Yes and AttackSeq-Procedure-No in each benchmark setting, where the bold value indicates the best performance of each column.
LLMs AttackSeq-Procedure-Yes AttackSeq-Procedure-No
@@ 3.3 Performance Comparison
Comparison on Contextual Information. Comparing the performance of LLMs across three benchmark settings in Table 2, we can observe that: In general, the Context setting consistently outperforms Zero-Shot and RAG settings across most benchmark tasks, with the advantage more pronounced in larger LLMs. Taking the Qwen-2.5 series as an example: performance shifts from zero-shot being optimal in Qwen-2.5-3B (0.4467 vs. 0.4614 vs. 0.3296) to context-setting being optimal in Qwen-2.5-32B (0.6195 vs. 0.5903 vs. 0.5154) in the AttackSeq-Tactic task, with Qwen-2.5-14B showing the transition in between. This phenomenon is reasonable as larger LLMs possess more extensive internal knowledge, and task-specific context further enhances their effectiveness and robustness within the specific domain. Moreover, both LLMs and LRMs consistently fail to reach optimal performance under the RAG-empowered setting. This indicates that naive retrieval integration may introduce additional noise instead of enhancing results, underscoring the requirement for more advanced retrieval-augmented approaches. We further investigate its limitation in Section 3.5.2.
3.4 Robustness Analysis
Figure 4: Parameter sensitivity analysis on (a) Temperature and (b) Max Output Tokens in AttackSeq-Tactic under the zero-shot setting.
3.4.1 Parameter Sensitivity Analysis
@@
Regarding the parameter sensitivity, we investigate the impact of temperature and maximum output tokens on LLMs and illustrate them in Figure 4(a) and Figure 4(b) respectively. Firstly, we observe that increasing the temperature from 0 to 1 causes a sharp performance drop in smaller LLMs, while larger LLMs remain relatively unaffected in attack sequence analysis. This may be because smaller LLMs lack discriminative power and oscillate among suboptimal answers, whereas larger ones generate more stable logits that preserve correct outputs even under smoothing, while aligning with the observations of previous work in general domain (Renze, 2024; Li et al., 2025). On the other hand, we also releaze that increasing the token budget yields stable performance and output length for LLMs, whereas LRMs achieve significant gains in both performance and output tokens. Specifically, when the Max Output Tokens increase from 1,024 to 4,096, R1 (LLaMA-8B) and R1 (Qwen-32B) improve accuracy by 13.29% and 16.74%, with average output tokens increasing by 37.35% and 43.28%, respectively. However, when the token budget is further increased to 8,192, LRMs exhibit diminishing returns: average output tokens increase by 13.27% and 9.83%, while accuracy improves only by 0.29% and 0.46%. This highlights the importance of carefully tuning the maximum output tokens parameter to optimize performance in LRMs while considering the associated costs incurred (Wang et al., 2024).
3.4.2 Computational Complexity Analysis
Figure 5: Computational complexity analysis of seven LLMs and four LRMs in AttackSeq-Tactic under the regular setting. The size of bubble represents inference time, where zigzag lines denote LLMs and cross hatch lines indicate LRMs.
@@ 3.4.2 Computational Complexity Analysis
As illustrated in Figure 5, we extensively compare the performance, model size and inference cost of several open-source LLMs and LRMs in AttackSeq-Tactic under the context setting. We can observe the following findings: (1) The performance trend of LLMs follows scaling law in our AttackSeqBench (Kaplan et al., 2020), particularly among LLMs within the same series; (2) Comparing the performance trend of LLMs with different thinking types, LRMs can not consistently outperform LLMs, yet they require significantly more inference time. This phenomenon is more pronounced when comparing LLMs and LRMs of the same model size, where distilled LRMs perform significantly worse than their base LLMs. This finding on inference time are consistent with existing studies in the general domain, whereas the performance results show a starkly contrasting trend (DeepSeek-AI, 2025). We further investigate this issue in Section 3.5.1.
@@
Table 4: A sample of reasoning path analysis between Qwen-2.5-32B (LLM) and R1(Qwen-2.5-32B) (LRM) on AttackSeq-Procedure-No under the zero-shot setting. We utilize different color to differentiate the diverse phases within a given attack sequence.
@@
Knowledge Retrieval: 1. Retrieving the Final Payload: This is a common step in many malware attacks where the payload is fetched from a command and control (C2) server or a malicious domain. 2. Generating Network Traffic: Interacting with the Transport Device Interface to generate network traffic could be a technique used to exfiltrate data or to communicate with a C2 server. However, this step seems out of place if the primary goal is to redirect the entry point of a driver. 3. Redirecting the Entry Point: This is a sophisticated technique used in firmware or UEFI-based attacks to gain persistence and control over the system at a very low level.
@@
(Transitional Word), but the question says ”only before redirecting.” So, does the TDI interaction happen only before the redirection, meaning it’s a one-time thing? Or is it part of the ongoing process? But wait (Transitional Word), could the redirection happen before the network traffic? (Questioning Alternatives) I think I’m leaning towards yes, so the answer is A: Yes. (Incorrect Conclusion)
3.5 In-depth Analysis
3.5.1 Reasoning Path Analysis
@@
To explore the reason why existing LRMs perform worser than LLMs in attack sequence understanding, we provide a sample of reasoning path between Qwen-2.5-32B (LLM) and R1 (Qwen-2.5-32B) (LRM) on AttackSeq-Procedure-No under the Zero-shot setting in Table 4. We observe that both LLM and LRM can decompose the attack sequence into granular TTPs and retrieve relevant knowledge. Here, LLM successfully aligns the retrieved knowledge with the logical order of TTPs, thereby recognizing that generating network traffic before redirecting the entry point contradicts the traditional attack sequence. It shows that LLM may rely on more direct sequence-matching between retrieved knowledge and procedural logic, enabling them to avoid unnecessary reasoning detours. In contrast, despite demonstrating reflective reasoning steps, LRM misinterprets the temporal constraint (“only before”) and overemphasizes the plausibility of the TDI interaction. This overthinking within LRMs are also more prone to construct redundant reasoning loops and further incur reasoning misalignment, which may amplify minor misunderstandings into incorrect conclusions.
@@
Figure 6: Error distribution of randomly-sampled 100 incorrect responses from GPT-4o in AttackSeq-Technique under the RAG-empowered setting.
3.5.2 Effectiveness of RAG Strategies
@@
To investigate why LLMs underperform in the RAG-empowered setting of our AttackSeqBench, we collect a candidate set where GPT-4o answers correctly in the zero-shot setting but fails under this setting. Specifically, we randomly sample 100 incorrect responses in AttackSeq-Technique, and classify them into four categories as shown in Figure 6. These four categories are: (1) Factual Error, meaning that LLM’s prediction contradicts the ground truth despite the correct retrieved content; (2) Over-reliance (Xia et al., 2024), meaning that LLM excessively refers to the retrieved content and fails to synthesize the attack sequence in the given question; (3) Irrelevant Retrieved TTP, which refers to incorrect predictions due to irrelevant retrieval to the given question; (4) Incorrect Answer Format, which refers to LLM’s failure to follow the output format specified in the prompt template.
@@
Our analysis reveals that 59% of errors stem from Factual Error, where the primary cause is the model’s failure to effectively integrate retrieved evidence into the reasoning chain. Rather than enhancing the inference process, the retrieved knowledge functions as noise to distort the output distribution, thereby inducing faulty reasoning and incorrect answers. Moreover, around 32% of the errors occur because LLMs treat retrieved knowledge as the absolute authority without validating them against the question intent or their internal knowledge. Consequently, the model often relies solely on correct but incomplete retrieval chunks, which leads to faulty results. Within the ATT&CK KB, the nuances of TTP descriptions introduce several overlaps and ambiguities, which account for 8% of cases where the embedding model to retrieve incorrect tactics and techniques. For example, technique T1574 – Hijack Execution Flow  [^2] https://attack.mitre.org/techniques/T1574/ is associated with three distinct tactics (i.e., Persistence, Privilege Escalation, and Defense Evasion), leading the model to misinterpret attack sequences in the given question. Enhancing the integration of retrieved knowledge with question intent and internal knowledge in RAG scenario, or investigating embedding methods capable of capturing fine-grained TTP semantics, holds promise for improving the effectiveness of attack sequence analysis.
4 Related Work
Automating CTI Report Understanding. With the increasing demands of cybersecurity operations and the breakthrough of LLMs, researchers have progressively explored their applicability within CRU (Zhang et al., 2024a). For instance, prior works have showcased the remarkable capabilities of LLMs in interpreting TTPs from the ATT&CK KB, surpassing the performance of some of the fine-tuned LMs with cybersecurity data (Fayyazi et al., 2024). Meanwhile, another line of work proposes LLM-driven threat intelligence Knowledge Graph construction frameworks, which utilize the threat-related entities and relations to describe CTI reports in a structural manner (Huang & Xiao, 2024; Cheng et al., 2024). However, the extent to which LLMs can understand and reason about the precise relations between adversary behavioral sequences described in CTI reports remains largely under-explored. In our work, we perform a holistic evaluation on various pre-trained LLMs, LRMs and fine-tuned LLMs in attack sequence analysis, from deducing high-level tactics to detailed procedures described in CTI reports.
Benchmarking LLMs in Cybersecurity. Inspired by the remarkable open-world knowledge and complex inference ability within LLMs, various benchmarks have been proposed to evaluate its general capabilities in language understanding (Hendrycks et al., 2021b), math reasoning (Cobbe et al., 2021), code generation (Chen et al., 2021). Regarding the cybersecurity domain, researchers start to benchmark the abilities of LLMs under such specialized setting, such as ethical hacking and compliance (Liu, 2023; Tihanyi et al., 2024; Garza et al., 2023). Targeting CRU-related tasks, SEvenLLM (Ji et al., 2024) explores the abilities of LLMs in threat-related entities extraction and summarizing reports from security vendors. SecBench (Jing et al., 2025) evaluates the knowledge retention and logical reasoning abilities of existing pre-trained LLMs from multiple languages and dimensions. Meanwhile, CTIBench (Alam et al., 2024) introduces five benchmark tasks to explore the threat entity attribution and cause-tracing abilities of LLMs within the security context.
5 Conclusion

### v3

@@ 4.1. Benchmark Settings
Zero-Shot Setting. Motivated by the significant Zero-Shot reasoning ability of LLMs across several downstream tasks (Hou et al., 2024; Kojima et al., 2022), we directly utilize the system prompt and Q&A pairs to evaluate LLMs’ performance on three tasks based on their inherent knowledge.
Figure 3. Overview of benchmark settings that exhibit varying levels of contextual information given to LLM.
Context Setting. Considering the existing context-aware work (Ma et al., 2023; Jin et al., 2024), we also organize context setting to evaluate LLMs’ Domain-Specific Epistemic Scalability at the semantic level. Here, we remove the corresponding summarized text $o_{k}$ of the ground truth tactic $t_{k}$ from the CTI outline $O$ to construct the masked CTI outline $O_{m}$, where $O_{m}\!=\!O\setminus o_{k}$. Afterwards, the LLM will be instructed to use the corresponding $O_{m}$ to answer the question, highlighting its potential to perform abductive reasoning to determine the most plausible TTP in attack sequence.
RAG-empowered Setting. Previous studies have demonstrated the Retrieval Augmented Generation (RAG) can significantly enhance the reliability of LLMs and mitigate hallucinations (Zhang et al., 2025a). Here, we also design the RAG-empowered setting to evaluate the Domain-Specific Epistemic Scalability of LLMs at the representation level. This leverages the LLMs’ in-context learning ability to learn the associations between the entity in the question’s body and the relevant TTPs, thereby decomposing the problem and eliciting its stronger reasoning ability (Wu et al., 2022).
4.2. Implementation Details
To investigate the CRU ability of existing models, we evaluate seven LLMs (i.e., LLaMa3.1-8B (Grattafiori et al., 2024), ChatGLM4-9B (GLM et al., 2024), Qwen2.5-3B, Qwen2.5-14B, Qwen2.5-32B (QwenTeam, 2024), LLaMa3.3-70B (Grattafiori et al., 2024), and GPT-4o (OpenAI, 2024a) ) and five LRMs (i.e., R1 (LLaMa3.1-8B), R1 (Qwen2.5-14B), R1 (Qwen2.5-32B) (DeepSeek-AI, 2025), QWQ-32B (Team, 2024) and GPT-o3-mini (OpenAI, 2025)) on AttackSeqBench. We use four post-training strategies (i.e., SFT (Zhang et al., 2023), RD (Huang et al., 2024), RLIF (Zhao et al., 2025) and RLVR (DeepSeek-AI, 2025)) to embed cybersecurity knowledge into LLMs to evaluate the Domain-Specific Epistemic Scalability of AttackSeqBench (cf. Appendix A.4). Here, we measure the performance with accuracy Acc=n/MAcc\!=\!n/M, where $n$ is the correctly-answered number of questions and $M$ is the total number.  [^2] We introduce the complete implementation details in Appendix A.5 and A.6.
Table 2. Performance comparison of various LLMs, LRMs and post-training strategies across three benchmark tasks and settings. Bold and underlined denote best and second-best in each column.
Versions Models AttackSeqBench-Tactic AttackSeqBench-Technique AttackSeqBench-Procedure
@@ 4.2. Implementation Details
4.3. Performance Comparison
Comparison between diverse groups of models. As shown in Table 2, we notice that: Although LLMs generally follow the scaling laws in our AttackSeqBench, none of them consistently outperforms the others, and the optimal LLM varies diverse tasks. For instance, the best-performing models under the Zero-Shot setting across three benchmark tasks are Qwen2.5-32B, Qwen2.5-32B, and GPT-o3-mini, respectively. This suggests that current models may not possess explicit security-specific knowledge, as relevant information in pre-training corpus is likely overshadowed by general-domain content. Moreover, most models consistently perform worst in AttackSeqBench-Tactic compared to other two tasks, mirroring the human evaluation results in Section 3.3 and underscoring the common challenge faced by both human experts and general LLMs in tactical inference.
Furthermore, we can observe that compared to the Zero-Shot setting, all models exhibit substantial performance gains on AttackSeqBench-Procedure-No under the context setting, indicating the importance of contextual information in identifying highly implausible actions within attack sequences. As defined in Appendix A.3, AttackSeqBench-Procedure-No is inherently more complex and reasoning-demanding than AttackSeqBench-Procedure-Yes, as it requires models to overcome the helpful-only bias and explicitly answer ‘No’ to disprove the plausibility of procedures occurring within the attack sequence. This explains why LRMs with stronger reasoning ability outperform in AttackSeqBench-Procedure-No compared to other tasks, underscoring the benchmark’s emphasis on Reasoning Scalability. Finally, most post-training strategies substantially improve the performance of its LLM backbone, particularly in Zero-Shot settings that rely solely on internal knowledge. However, their performance still lags behind instructive LLMs equipped with task-adapted prompts. This highlights a promising direction: Designing a specialized post-training strategy to embed security-related knowledge into existing LLMs, thereby advancing the development of domain-specific models for cybersecurity domain.
Table 3. Performance comparison of four representative LLMs and LRMs under each setting, where the bold value indicates the best performance of each column.
LLMs AttackSeqBench-Procedure-Yes AttackSeqBench-Procedure-No
@@ 4.3. Performance Comparison
Comparison on Contextual Information. Comparing the performance of LLMs across three benchmark settings in Table 2, we can observe that: In general, the Context setting consistently outperforms Zero-Shot and RAG settings across most benchmark tasks, with the advantage more pronounced in larger LLMs. Taking the Qwen2.5 series as an example: performance shifts from Zero-Shot being optimal in Qwen2.5-3B (0.4467 vs. 0.4614 vs. 0.3296) to context-setting being optimal in Qwen2.5-32B (0.6195 vs. 0.5903 vs. 0.5154) in AttackSeqBench-Tactic, with Qwen2.5-14B showing the transition in between. This phenomenon is reasonable as larger LLMs possess more extensive internal knowledge, and task-specific context further enhances their effectiveness and robustness within the specific domain. Moreover, both LLMs and LRMs consistently fail to reach optimal performance under the RAG-empowered setting. This indicates that naive retrieval integration may introduce additional noise instead of enhancing results, underscoring the requirement for more advanced retrieval-augmented approaches. We further investigate its limitation in Section 4.5.2.
$4.$ Robustness Analysis
4.4.1. Parameter Sensitivity Analysis
@@
Regarding the parameter sensitivity, we investigate the impact of temperature and maximum output tokens on LLMs and LRMs and illustrate them in Figure 4 (a) and Figure 4 (b) respectively. Firstly, we observe that increasing the temperature from 0 to 1 causes a sharp performance drop in smaller LLMs, while larger LLMs remain relatively unaffected in attack sequence analysis. This may be because smaller LLMs lack discriminative power and oscillate among suboptimal answers, whereas larger ones generate more stable logits that preserve correct outputs even under smoothing, while aligning with the observations of previous work in general domain (Renze, 2024; Li et al., 2025). On the other hand, we also releaze that increasing the token budget yields stable performance and output length for LLMs, whereas LRMs achieve significant gains in both performance and output tokens. Specifically, when the Max Output Tokens increase from 1,024 to 4,096, R1 (LLaMA-8B) and R1 (Qwen-32B) improve accuracy by 13.29% and 16.74%, with average output tokens increasing by 37.35% and 43.28%, respectively. However, when the token budget is further increased to 8,192, LRMs exhibit diminishing returns: average output tokens increase by 13.27% and 9.83%, while accuracy improves only by 0.29% and 0.46%. This highlights the importance of carefully tuning the maximum output tokens parameter to optimize performance in LRMs while considering the associated costs incurred (Wang et al., 2024).
Figure 4. Parameter sensitivity analysis on (a) Temperature and (b) Max Output Tokens in AttackSeqBench-Tactic.
4.4.2. Computational Complexity Analysis
@@
As illustrated in Figure 5, we extensively compare the performance, model size and inference cost of several open-source LLMs and LRMs in AttackSeqBench-Tactic under the context setting. We can observe the following findings: (1) The performance trend of LLMs follows scaling law in our AttackSeqBench (Kaplan et al., 2020), particularly among LLMs within the same series. (2) Comparing the performance trend of LLMs with different thinking types, LRMs can not consistently outperform LLMs, yet they require significantly more inference time. This phenomenon is more pronounced when comparing LLMs and LRMs of the same model size, where distilled LRMs perform significantly worse than their base LLMs. This finding on inference time are consistent with existing studies in the general domain, whereas the performance results show a starkly contrasting trend (DeepSeek-AI, 2025). We further investigate this issue in Section 4.5.1.
@@
Table 4. A sample of reasoning path analysis between Qwen2.5-32B and R1 (Qwen2.5-32B) on AttackSeqBench-Procedure under the Zero-Shot setting. We utilize different color to differentiate the diverse phases within a given attack sequence.
@@
Knowledge Retrieval: 1. Retrieving the Final Payload: This is a common step in many malware attacks where the payload is fetched from a command and control (C2) server or a malicious domain. 2. Generating Network Traffic: Interacting with the Transport Device Interface to generate network traffic could be a technique used to exfiltrate data or communicate with a C2 server. However, this step seems out of place if the goal is to redirect the entry point of a driver. 3. Redirecting the Entry Point: This is a sophisticated technique used in firmware or UEFI-based attacks to gain persistence and control over the system at a very low level.
@@
(Transitional Word), but the question says "only before redirecting." So, does the TDI interaction happen only before the redirection, meaning it’s a one-time thing? Or is it part of the ongoing process? But wait (Transitional Word), could the redirection happen before the network traffic? (Questioning Alternatives) I think I’m leaning towards yes, so the answer is A: Yes. (Incorrect Conclusion)
4.5. In-depth Analysis
4.5.1. Reasoning Path Analysis
@@
To explore the reason why LRMs perform worser than LLMs in our AttackSeqBench, we provide a sample of reasoning path between Qwen2.5-32B and R1 (Qwen2.5-32B) in Table 4. We observe that both LLM and LRM can decompose the attack sequence into granular TTPs and retrieve relevant knowledge. Here, LLM successfully aligns the retrieved knowledge with the logical order of TTPs, thereby recognizing that generating network traffic before redirecting the entry point contradicts the traditional attack sequence. It shows that LLM may rely on more direct sequence-matching between retrieved knowledge and procedural logic, enabling them to avoid unnecessary reasoning detours. In contrast, despite demonstrating reflective reasoning steps, LRM misinterprets the temporal constraint (“only before”) and overemphasizes the plausibility of the TDI interaction. This overthinking within LRMs are also more prone to construct redundant reasoning loops and further incur reasoning misalignment, which may amplify minor misunderstandings into incorrect conclusions.
Figure 5. Computational complexity analysis of seven LLMs and four LRMs in AttackSeqBench-Tactic under the regular setting. The size of bubble represents inference time, where zigzag lines denote LLMs and cross hatch lines indicate LRMs.
@@
Figure 6. Error distribution of randomly-sampled 100 incorrect responses from GPT-4o in AttackSeqBench-Technique under the RAG-empowered setting.
4.5.2. Effectiveness of RAG Strategies
@@
To investigate why LLMs underperform in the RAG-empowered setting of our AttackSeqBench, we collect some candidates where GPT-4o answers correctly in Zero-Shot setting but fails under this setting. Specifically, we randomly sample 100 incorrect responses in AttackSeqbench-Technique, and classify them into four categories in Figure 6. These four categories are: (1) Factual Error, meaning that LLM’s prediction contradicts the ground truth despite the correct retrieved content; (2) Over-reliance (Xia et al., 2024), meaning that LLM excessively refers to the retrieved content and fails to synthesize the attack sequence in the given question; (3) Irrelevant Retrieved TTP, which refers to incorrect predictions due to irrelevant retrieval to the given question; (4) Incorrect Answer Format, which refers to LLM’s failure to follow the output format specified in the prompt template.
@@
Figure 6 reveals that 59% of errors stem from Factual Error, where the primary cause is the model’s failure to effectively integrate retrieved evidence into the reasoning chain. Rather than enhancing the inference process, the retrieved knowledge functions as noise to distort the output distribution, thereby inducing faulty reasoning and incorrect answers. Moreover, 32% of errors occur because LLMs treat retrieved knowledge as the absolute authority without validating them against the question intent. Consequently, the model often relies solely on correct but incomplete retrieved chunks, which leads to faulty results. Within the ATT&CK KB, the nuances of TTP descriptions introduce several overlaps and ambiguities, which account for 8% of cases where the embedding model to retrieve incorrect tactics and techniques. For example, technique T1574 – Hijack Execution Flow  [^3] https://attack.mitre.org/techniques/T1574/ is associated with three distinct tactics (i.e., Persistence, Privilege Escalation, and Defense Evasion), leading the model to misinterpret the given attack sequences. Enhancing the integration of retrieved knowledge with question intent, identifying SAE-derived feature to generate answers that aligned with retrieved evidence, or tuning the embedding models to capture more accurate TTP semantics, holds promise for improving the effectiveness of RAG in attack sequence understanding.
5. Conclusion